OpenStudy (anonymous):

4 + 2(x − 1) = 5x − (x + 10)

OpenStudy (anonymous):

First the equation is 4 + 2(x − 1) = 5x + (−1)(x + 10), Distribute the 2 and the -1 to get 4 + 2x - 2 = 5x - x - 10

OpenStudy (anonymous):

Add like terms you will get 2x + 2 = 4x - 10, subtract 2 from both sides and get: 2x = 4x - 12 then subtract 4x from both sides and get -2x = - 12 and divide by -2

OpenStudy (anonymous):

You will get x = 6

OpenStudy (anonymous):

Thank you @TheMusician your very helpful :)

OpenStudy (anonymous):

No problem. TY for the compliment. :D

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Love your pic also!